OpenStudy (anonymous):

write an example of an algebraic expression that always has the same value of the variable???

OpenStudy (anonymous):

always has same value reguardless value of variable

Directrix (directrix):

Have you studied the reflexive property?

OpenStudy (anonymous):

yeah last year in geometry im in algebra 2. would it be eggs - eggs?? for the answer

Directrix (directrix):

I am not following how eggs come into the solution. >eggs - eggs

OpenStudy (anonymous):

it doesn't say that in question. my friend thought that might be the answer. I wasn't sure

Directrix (directrix):

I was thinking along the lines of x = x or maybe 2*x = 0. The expression must have the same value as the variable. On 2*x = 0, x has to be 0. I don't know if that is what this question is about.

Directrix (directrix):

On 2*x = 0, the 2*x would have to be 0 if 2*x = 0.

OpenStudy (anonymous):

that doesn't make sense with my question I don't think??? your confusing

Directrix (directrix):

Well, I'm not sure I understand the question so I'll move out of the way and perhaps someone else can see the problem clearly.

OpenStudy (ybarrap):

A simple algebraic expression that always has the same value might be something like x=6. The variable x always has the same value: 6.

OpenStudy (anonymous):

If an expression has the exact same value regardless of the variable, then it would follow that the variable itself shouldn't be in the expression itself. Either that or the expression is one that cancels the variable to 0. So, such an example would be y = 5 or y = 0. If you were to plug in any x at all, you'd find that there isn't an x in the equation, so the value would always be y = 5 or y = 0.

OpenStudy (anonymous):

thanks!! ybarrap

OpenStudy (anonymous):

thanks?

OpenStudy (skullpatrol):

If a is a real number, then $$\Huge a=a$$

OpenStudy (anonymous):

thanks skullpatrol I understood that

OpenStudy (anonymous):

mind answering more questions? and is the profile pic what u really look like or a google pic?

OpenStudy (anonymous):

In a soccer tournament, teams receive 6 points for a winning game, 3 points for a trying game, and 1 point for each goal they score. what algebraic expression models the total number of points that a soccer team receives in a tournament?????? suppose one team wins two games and ties 1 game, scoring a total of 5 goals. how many points does the team receive????